OpenStudy (anonymous):

At what height from surface of earth, acceleration due to gravity becomes g/2 ?

OpenStudy (fifciol):

\[\frac{ g }{ 2}=\frac{ GM }{2R ^{2} }= \frac{ GM }{(R+h) ^{2} }\rightarrow h=R(\sqrt{2}-1)= 2639 km\]

OpenStudy (anonymous):

2.6km aproximately